Please Help Me Solve This Quadratic Equation Application:
The Border of uniform width around a pool has outer dimensions of 26 feet by 66 feet. Find the width of the border if the pool has an area of 1200 square feet

Answer by checkley77(12844)

(26-2x)(66-2x)=1,200
1,716-132x-52x+4x^2=1,200
4x^2-184x+1,716-1,200=0
4x^2-184x+516=0
4(x^2-46x+129)=0
4(x-43)(x-3)=0
x-3=0
x=3 ft. is the width of the border around the pool.
Proof:
(26-2*3)(66-2*3)=1,200
(26-6)(66-6)=1,200
20*60=1,200
1,200=1,200